Ten cups of a restaurants' house Italian dressing is made by blending olive oil costing $1.50 per cup with vinegar that costs $.25 per cup. How many cups of each are used if the cost of the blend is $.50 per cup?
Answer by funmath(2933)

Ten cups of a restaurants' house Italian dressing is made by blending olive oil costing $1.50 per cup with vinegar that costs $.25 per cup. How many cups of each are used if the cost of the blend is $.50 per cup?
Let the amount of oil be: x
The the rest of the 10 cups is the amount of vinegar: 10-x
The total mixture amount is: 10 cups
The cost of oil is: 1.50x
The cost of vinegar is: .25(10-x)
The cost of the mixture is: .50(10)=5
The problem to solve is:
1.50x+.25(10-x)=5
1.50x+2.5-.25x=5
1.25x+2.5=5
1.25x+2.5-2.5=5-2.5
1.25x=2.5
1.25x/1.25=2.5/1.25
x=250/125
x=2
The amount of oil is: x=2 cups
The amount of vinegar is: 10-x=10-2=8 cups
Sanity check, does this make sense?
1.50(2)+.25(8)=.5(10)
3+2=5
5=5 yes.
Happy Calculating!!!!